OpenStudy (anonymous):

At the end of her aunt's story in “The Story-Teller," the older girl asks whether the girl in the story would have been rescued even if she had not been good. What does she seem to think is the theme of her aunt's story? What detail in the aunt's story leads her to think so?

OpenStudy (anonymous):

she thought \(\sum_{n=1}^{\infty}\frac{1}{n^2}=\frac{\pi^2}{6}\)